OpenStudy (anonymous):

list all x intercepts for y= -3sin(1/2x + 1/5 pi) on the interval [-2/5pi,4pi]

OpenStudy (anonymous):

substitute y = 0

OpenStudy (anonymous):

did that and got stuck :-)

OpenStudy (anonymous):

good idea satellite - i'm getting a nice cold beer now!

OpenStudy (anonymous):

sin (x/2 + pi/5)=0 what are all the possible values of theta for sin theta = 0?

OpenStudy (anonymous):

how come you removed the -3 from sin?

OpenStudy (anonymous):

y = 0 = -3sin(1/2x + 1/5 pi) -3(1/2x + 1/5 pi) = 0 divide by 3 on both sides sin(1/2x + 1/5 pi)=0

OpenStudy (anonymous):

I got x= 5.03, -1.26, 11.31

OpenStudy (anonymous):

ok, thanks! loophole can you please show me how

OpenStudy (anonymous):

I am really struggling with this one

OpenStudy (anonymous):

I used a program called Geogebra, so I didn't get it algebraically sorry, but you can check your answers with that. I'll try it now though.

myininaya (myininaya):

i got 18pi/5, 8pi/5,-2pi/5

OpenStudy (anonymous):

myininaya our answers are equal

OpenStudy (anonymous):

you should get 5 answers for the interval specified.

OpenStudy (anonymous):

that is one of the answers: i am more concerned with understanding why - thanks everyone so much!

myininaya (myininaya):

if you can read it this is what dhatraditya was trying to explain above

OpenStudy (anonymous):

can;t be 5 answers; the max in the range of answers is 3

OpenStudy (anonymous):

there are only 3

OpenStudy (anonymous):

how is that? there will be one at 0,pi,2pi,3pi and 4pi

myininaya (myininaya):

when i did 3pi dhatraditya i got someing over 4pi for the answer

OpenStudy (anonymous):

the range part of the question

OpenStudy (anonymous):

wait, the interval is specified for the angles that the sin value can take or the resultant Y can take?

OpenStudy (anonymous):

thank you so much myninaya! that is perfect!!!! can some one please recommend a decent place for me to brush up on unit circle trig

myininaya (myininaya):

x is between -2pi/5 and4pi when i did theta =4pi , x was bigger than 4pi

myininaya (myininaya):

thetha=3pi*

OpenStudy (anonymous):

oh I see. silly me,.

myininaya (myininaya):

so we have only 3 solutions

OpenStudy (anonymous):

how would the answer change if sin were replaced by cos?

myininaya (myininaya):

i eat dhartrad got this right?

OpenStudy (anonymous):

not much, your angles would become pi/2, 3pi/2 and so on

OpenStudy (anonymous):

@myininaya. yup, got it.

OpenStudy (anonymous):

you guys are fantastic! Im going to post it now with cos :-)